OpenStudy (anonymous):

Simplify the expression 6√2 over √3.

OpenStudy (anonymous):

\[[6\sqrt{2}/\sqrt{3}] [\sqrt{3}/\sqrt{3}]\]

OpenStudy (anonymous):

6\[\sqrt{2*3}\] over \[\sqrt{9}\]

OpenStudy (anonymous):

6 times \[\sqrt{2*3}\] is that

OpenStudy (anonymous):

then 6 time sqrt(6) over 3 = 2 sqrt(6)